OpenStudy (anonymous):

Solve for x: |x + 2| + 16 = 14 x = −32 and x = −4 x = −4 and x = 0 x = 0 and x = 28 No solution

OpenStudy (anonymous):

\[\left| x+2 \right|=14-16=-2\] absolute value cannot be negative. Hence no solution.